OpenStudy (anonymous):

Solve the equation for x: 2x - 16y = 10. x = 10 + 8y x = 5 - 16y x = 5 + 8y x = 5 + 16y

OpenStudy (anonymous):

how you solve for this?

OpenStudy (anonymous):

2x-16y=10 2x=10+16y x=(10+16y)/2 x=5+8y (c)